A 3m long diving board weighing 50 N is supported at one end and at a pivot 1 m from the end. At the other end stands a diver weighing 100 N. What are the forces on the board from the support and the pivot?
Answer #1

here,

length of diving board , l = 3 m

weight of board , W1 = 50 N

weight of diver , W2 = 100 N

let the forces on the board from the support and the pivot be Fs and Fp

taking moment of force about the suport

Fp * 1 - W1 * l/2 - W2 * l = 0

Fp - 50 * 3/2 - 100 * 3 = 0

Fp = 375 N

taking moment of force about the pivot

- Fs * 1 = W1 * ( l/2 - 1) + W2 * (l - 1)

- Fs = 50 * 0.5 + 100 * 2

Fs = - 225 N

the forces on the board from the support and the pivot is (225 downwards) and (375 upwards)